OpenStudy (yb1996):

Find the second derivative of (2x+1)^2/(sqrt (x))

OpenStudy (amistre64):

quotients tend to be simpler if you make them into a product

OpenStudy (yb1996):

Ok, can you please show me how that's done? I've tried to solve this problem about 5 times and keep on getting the wrong answer.

OpenStudy (freckles):

have you found the first derivative yet?

OpenStudy (amistre64):

f = uv f' = u'v + uv' f'' = u''v + u'v' + u'v' + uv''

OpenStudy (amistre64):

let u = the top part, and v = the bottom, but recall that 1/x = x^(-1)

OpenStudy (amistre64):

these notices are not working properly at the moment

OpenStudy (freckles):

\[f=uv \\ f'=u'v+uv' \\f''=u''v+2u'v'+uv'' \\ f'''=u'''v+u''v'+2u''v'+2u'v''+u'v''+uv''' \\=u^{(3)}v+3u^{(2)}v^{(1)}+3u^{(1)}v^{(2)}+uv^{(3)}\] pascal's triangle :) anyways this wasn't all needed for your thingy your problem whatever

OpenStudy (yb1996):

Ok, thanks guys. I'll give it a shot.